OpenStudy (knudsona):

I'm in virtual school. And I keep getting this question wrong. Could anyone answer. and then explain how they got that answer. Appreciated. :) Evaluate the following expression using the values given: Find -3a2 - b3 + 3c2 - 2b3 if a = 2, b = -1, and c = 3.

OpenStudy (justnick09):

okay so plug in the equation

OpenStudy (knudsona):

I did. and when i submitted my answer they said it was incorrect?

OpenStudy (anonymous):

the answer is 18

OpenStudy (justnick09):

so it becomes\[-6(2)-3+9(2)+2(3)\]

OpenStudy (anonymous):

hope this helps

OpenStudy (anonymous):

\[\large -3a^2 - b^3 + 3c^2 - 2b^3\] \[=\large -3a^2 - 3b^3 + 3c^2\]is a start

OpenStudy (justnick09):

\[-12-3+18\]

OpenStudy (anonymous):

is it right? @knudsona

OpenStudy (justnick09):

the final answer is 3

OpenStudy (anonymous):

\[=\large -3a^2 - 3b^3 + 3c^2\] \[=\large -3\times ^2 - 3\times(-1)^3 + 3\times 3^2\]

OpenStudy (knudsona):

I feel insanely stupid right now.

OpenStudy (justnick09):

look at what i gave u and try it

OpenStudy (anonymous):

why?

OpenStudy (anonymous):

\[=\large -3\times2 ^2 - 3\times(-1)^3 + 3\times 3^2\]

OpenStudy (anonymous):

@knudsona Why do you feel stupid

OpenStudy (justnick09):

satelite explain urself i dont even get where your going

OpenStudy (knudsona):

@justnick09 your answer is incorrect.

OpenStudy (anonymous):

\[-3\times 4-3\times (-1)+3\times 9\] \[-12+3+27\]\[-9+27=18\]

OpenStudy (justnick09):

ohhhh those were in powers omg...

OpenStudy (justnick09):

use these for powers 7^2

OpenStudy (knudsona):

yeah sorry, i didnt know what t do for that. lol.

OpenStudy (justnick09):

@satellite73 is right

OpenStudy (knudsona):

thank you @satellite73 @theonlycool13 @justnick09

OpenStudy (anonymous):

hey satellite how did you get what you did i don't quite follow

OpenStudy (anonymous):

yo welcome